OpenStudy (anonymous):

ind_int xe^6x dx u=x du=dx dv=e^6x dx v=(1/6)e^6x

OpenStudy (anonymous):

what do you need help with?

OpenStudy (anonymous):

your substitutions are correct. where did you get stuck?

OpenStudy (anonymous):

\[the final ans is (x/6)e ^{6x}-(1/36)e ^{6x}+c....don't get the (1/36) part?? how is \int\limits_{}^{}(1/6)e ^{6x}=(1/36)?\]

OpenStudy (anonymous):

\[\int\limits_{}^{}(1/6)e ^{6x}=(1/36)e ^{6x}?\]

OpenStudy (anonymous):

From Paul's Online CalcII notes.

OpenStudy (anonymous):

i'll just solve it step by step.. \[uv - \int vdu\] \[x/6 (e^{6x}) - \int 1/6 (e^{6x})dx\] \[x/6(e^{6x}) - 1/6 \int e^{6x}dx\] the integral of e^(6x)dx is 1/6 e^(6x) \[x/6(e^{6x}) - 1/6 (e^{6x}/6)\] \[x/6(e^{6x} - 1/36 (e^{6x})\]

OpenStudy (anonymous):

It's the obvious points that elude me. Exactly why I'm not so great at math. THanks for your help!

OpenStudy (anonymous):

you're welcome :)